OpenStudy (anonymous):

what affect does k=5 have on the parent function y=1/x

OpenStudy (solomonzelman):

well in your case, "y=1/x" the k is 1. if you make it 5,it is going to make x less than when k=1.

OpenStudy (solomonzelman):

As in any inverse variation when you have y=k/x the greater the k the greater the ration is between y and x.

OpenStudy (solomonzelman):

I meant ratio not ration....

OpenStudy (anonymous):

so it would move down?

OpenStudy (solomonzelman):

y=k/x yx=k the greater the k the bigger the product of x and y is, but not necessarily I think....